### v1

@@ 1 Introduction
2 Related Works
The Stitching Property The concept of stitching, as discussed by (Ziebart et al., 2008), is a characteristic feature of TD-based RL algorithms, such as those described by (Mnih et al., 2013), (Lillicrap et al., 2015), (Fujimoto et al., 2018), and (Kostrikov et al., 2021), which employ Dynamic Programming techniques. This feature allows these algorithms to integrate data from different trajectories, thereby enhancing their effectiveness in managing complex tasks by leveraging historical data (Cheikhi and Russo, 2023). Recent works have shown that self-supervised learning also exhibits such property, particularly through methods like data augmentation (Shorten and Khoshgoftaar, 2019; Ghugare et al., 2024) . However, there remains a significant performance gap between these methods and algorithms that incorporate Dynamic Programming.
@@ 2 Related Works
Goal-conditioned Weighted Supervised Learning Goal-conditioned weighted supervised learning methods (GCWSL) (Liu et al., 2021; Yang et al., 2022; Ma et al., 2022; Hejna et al., 2023) are new outstanding goal-conditioned algorithms which enable agent to reach multiple goals in GCRL. These methods repeatedly perform imitation learning on self-collected, relabeled data without any RL updates. Furthermore, these methodologies undergo theoretical validation to confirm that self-supervised learning from relabeled experiences optimizes a lower bound on the GCRL objective. This validation guarantees that the learning process is both efficient and underpinned by a rigorous theoretical foundation.
@@ 3.1 Goal-conditioned Reinforcement Learning
Q-learning methods optimizes $\mathcal{J}(\pi)$ via the gradient computation:
@@ 3.2 Goal-conditioned Weighted Supervised Learning
where $\mathcal{B}_{r}$ denotes relabeled replay buffer, $g=\phi(s_{i})$ denotes the relabeled goals for $i\geq t$ and $f(A)$ is a function about advantage $A:=A(s_{t},a_{t},g)$. The weighted function $f(A)$ exists various forms in GCWSL methods (Ghosh et al., 2021; Yang et al., 2022; Ma et al., 2022; Hejna et al., 2023). (See Appendix A.3 for the specific weight function meaning). Therefore GCWSL includes typical two process, acquiring sub-trajectories corresponding to state-goal (i.e, $(s,g)$) pairs and imitating them. In the process of imitation, GCWSL first train the related specific weighted function, and then extract the policy with the Eq. 4.
@@
Theorem 1
@@
$\ \pi_{k+1}=\underset{\pi\in\Pi}{\operatorname*{\arg\max}}\mathbb{E}_{(s,g,s^{\prime})\sim\mathcal{B}_{r},a_{t}\sim\pi(s,g)}\left[A^{\pi_{k}}(s,a,g)\right],$
[…]
where $A^{\pi_{k}}=Q^{\pi_{k}}(s,a,g)-V^{\pi_{k}}(s,g)$, $\mathcal{D}_{\mathrm{KL}}$ is the KL-divergence and $g$ is the goals from relabeled data. See Appendix A.1 for detail proof. This theorem indicate GCWSL guarantees policy improvement when learning from the relabelled data.
@@ 4.1 Trajectory Stitching Example
Consider an example illustrated in Fig. 1, where the goals $g\in\left[g^{a},g^{b}\right]$ are reachable from the states $s\in\left[s_{0}^{a},s_{0}^{b},s_{t}\right]$ although not explicitly present in the off-policy replay buffer. During training, GCWSL is updated with (state, goal) pairs sampled from the same trajectory, such as $(s_{0}^{a},g^{a})$ and $(s_{0}^{b},g^{b})$. However, during testing, when unseen (state, goal) pairs such as $(s_{0}^{a},g^{b}),(s_{0}^{b},g^{a})$ originating from distinct trajectories are introduced, GCWSL cannot find the correct actions (proof can be found in Figure 1 and Lemma 4.1 of (Ghugare et al., 2024)). In contrast, TD-based RL methods such as Q-learning can output the correct actions for $(s_{0}^{a},g^{b}),(s_{0}^{b},g^{a})$ due to the Dynamic Programming component. For instance, consider a value estimation example, $V(s_{t},g)=r+V(s_{t+1},g)$. For the pair $(s_{t+1},g)$, provided that $g$ is accessible and the value $V(s_{t+1},g)$ is accurately approximated, Q-learning can back-propagate the state-goal (state-action-goal) values to preceding pairs. An example from the off-policy replay buffer illustrates the state $s_{t}$ as a "hub state" shared by two initial trajectories, from which all goals $g^{i}\sim[g^{a},g^{b}]$ are reachable. The values of these goal-conditioned states $Q(s_{t},a_{t}^{i},g^{i})$ can be computed and recursively back-propagated to $Q(s_{0},a_{0}^{i},g^{i}),i\in\left[a,b\right]$. Q-learning has the capability to propagate rewards through paths that are connected retrospectively. This allows Q-learning to determine the optimal actions for state-goal pairs that have not been encountered together previously, such as $(s_{0}^{a},g^{b})$ and $(s_{0}^{b},g^{a})$. By iteratively back-propagating value estimates through all potential trajectory combinations, Q-learning ensures that values are consistently updated, thereby facilitating the correct action selection. In summary, Q-learning can link (state, goal) pairs across different trajectories, a capability that GCWSL lacks.
@@ 4.2 Bounding KL-Constrained Values and Q-WSL
If we he replace the advantage function $A^{\pi}(s,a,g)$ in Eq. 1 by Q-function $Q^{\pi}(s,a,g)$ we have the policy improvement objective:
$\pi_{k+1}=\arg\max_{\pi}\mathbb{E}\left[Q^{\pi_{k}}(s,\pi_{k}(s,g),g)\right],$
$\;\int_{a}\pi(a|s,g)\boldsymbol{d}a=1,\;\mathbf{s.t.}~{}\operatorname{KL}(\pi\|\pi_{relabel})\leq\epsilon.$ (6)
@@
$\ \mathcal{L}(\lambda,\pi^{H})=\mathbb{E}_{a\sim\pi(\cdot|s,g)}\left[Q^{\pi}(s,a,g)\right]+$
$\lambda\mathbb{E}_{a,s,g\sim\mathcal{B}_{r}}\log(\pi(a|s,g)).$ (7)
For a deterministic policy $\pi$, it can be regarded as a Dirac-Delta function. Consequently, the constraint $\int_{a}\pi(a|s,g)\boldsymbol{d}a\\ =1$ is always satisfied. Therefore, Equation Eq. 4.2 simplifies to:
@@
Based on this discussion and the interoperation of these two objectives, we combine Eq. 4.2 with Eq. 1 to optimize the policy concurrently. This leads to our Q-WSL policy optimization objective for policy updates:
$\ \mathcal{J}_{Q-WSL}=\mathbb{E}_{(s,a,g)\sim\mathcal{B}_{r}}\big{[}\underbrace{Q^{\pi}(s,a,g}_{\text{Q-learning}})$
$+\eta\underbrace{f(A^{\pi}(s,a,g))\cdot\log\pi_{\theta}(s,a,g)}_{\text{Advantage-weighted Regression}}\big{]},$ (10)
@@ 4.3 Practice Algorithm
We adopt the DDPG+HER as our baseline framework following (Fang et al., 2019; Yang et al., 2021), therefore for a deterministic policy $\pi(s,g)$, the Eq. 20 and Eq. 8 are used during training procedure. Moreover, we follow WGCSL (Yang et al., 2022) to define $f(A)$ equal to $\gamma^{i-t}\exp_{clip}(A(s_{t},a_{t},g))\cdot\epsilon(A(s_{t},a_{t},g))$, where $A(s_{t},a_{t},g)=r(s_{t},a_{t},g)+\gamma V(s_{t+1},g)-V(s_{t},g)$ and the value function $V(s_{t},g)$ is estimated using Q-value function as $V(s_{t},g)=Q(s_{t},\pi(s_{t+1},g),g)$ Because in the experiment, we found that the effect of this form of $f(A)$ is better for online goal-conditioned RL, and other forms of $f(A)$ may be more friendly for offline settings. The final overview of Q-WSL is presented in Algorithm 2.
Compared with WGCSL Motivated by (Chane-Sane et al., 2021; Yang et al., 2022), we also compared with WGCSL and have the following boundary:
Theorem 2
Suppose function $\exp_{clip}(A(s_{t},a_{t},g)\cdot\epsilon(A(s_{t}$ $,a_{t},g))\geq 1,g=\phi(s_{i}),i\geq 1$ over the state-action-goal combination. Consider a finite-horizon discrete Markov Decision Process (MDP), a deterministic policy $\pi$ that selects actions with non-zero probability, and a sparse reward function $r(s_{t},a_{t},g)=\mathbb{1}\left[\phi(s_{i})=g\right]$, where $1\left[\phi(s_{i})=g\right]$ is an indicator function. Given trajectories $\tau=(s_{1},a_{1},\cdots,s_{T},a_{T})$ and a discount factor $\gamma\in(0,1]$, the following bounds hold:
@@
We defer the proof to Appendix A.1. This theorem indicates that under a reward of $r(s_{t},a_{t},g)=\mathbb{1}\left[\phi(s_{i})=g\right]$, the optimization objective of Q-WSL is at least better than that of WGCSL and GCSL.
Algorithm 2 Q-WSL
@@
13:         Update $\pi_{\theta}$ with maximize goal-conditioned policy optimization with Eq. 20 and Eq. 8
@@
4.4 Compared with Previous Goal-conditioned Methods
@@ 5 Experiments
Benchmarks We use some standard goal-conditioned RL research benchmarks (Plappert et al., 2018) including one task on the $Shadow-hand$ and four tasks on the $Fetch$ robot (See Fig. LABEL:fig:multi_goal_tasks).
Baselines In this section, we compare our method Q-WSL with prior works on various goal-conditioned policy learning optimization algorithms as show in Table 1 . We conducted a series of comparative experiments by implementing the baseline methods within the same framework as our proposed approach.
Detailed baselines algorithm implementation is shown in Appendix A.3. We provide a brief overview of our baseline algorithms as bellow:
@@
Architecture Setup We utilize the DDPG (Paul et al., 2018) in conjunction with HER as the foundational algorithm for goal-conditioned RL, against which we benchmark various baselines. Specifically, the critic architecture $Q(s,a,g)$ adheres to the Universal Value Function Approximators (UVFA) model (Schaul et al., 2015). The actor network generates actions parameterized by a Gaussian distribution. The actor and critic networks are structured with three hidden layers of multi-layer perceptrons (MLPs), each comprising 256 neurons and employing ReLU activations. The architectural setup can be denoted as [linear-ReLU]×3, followed by a linear output layer. We follow the optimal parameter settings for DDPG+HER as specified in (Plappert et al., 2018). All baseline hyperparameters are maintained consistently. Additional details are provided in Appendix A.5.
[…]
500}500\}500 }). The policy was trained for 1000 episodes per epoch. Upon completing each training epoch, the policy’s performance was measured by calculating the average success rate from 100 independent rollouts, each using randomly selected goals. These success rates were averaged across five seeds and plotted over the learning epochs, with the standard deviation illustrated as a shaded region on the graph.
@@
5.1 Performance Evaluation on Multi-goal Benchmark Results
@@ 5.2 Sample Efficiency
To evaluate the sample efficiency of Q-WSL compared to baseline methods, we analyze the number of training samples (i.e., $\left\langle{s,a,g^{\prime},g}\right\rangle$ tuples, $g^{\prime}$ is the relabeled goal) required to attains a specific mean success rate. For this comparison, we have selected competitive baselines that employ goal relabeling strategies in the HandReach environment.
The results are depicted in Fig. 6. In the HandReach environment, as shown in Fig. 6, Q-WSL requires only 1000 samples to attains the 0.1 mean success rate, demonstrating superior sample efficiency. In this case, Q-WSL is at least twice the sample efficiency of previous works.
@@
The reason why a state-of-the-art baseline DWSL is not listed here is that there is basically no success rate in the HandReach environment. Overall, Q-WSL enhances sample efficiency by an average factor of seven (6.8) compared to the baseline methods.
@@ 5.3 Robust to Environmental Stochasticity
As we see in Fig. 7, Q-WSL is the most robust to stochasticity in the FetchPush environment, also outperforming baseline algorithms in terms of mean success rate under various noise levels. WGCSL exhibits minimal sensitivity to variations in noise levels, whereas DDPG+HER
@@
is more sensitive, and GoFar is the most affected by noise fluctuations. Despite DWSL’s insensitivity to noise, its overall performance remains suboptimal. We analyze that the assumption of deterministic dynamics in the self-supervised learning methods such as WGCSL, GoFar, and DWSL may lead to overly optimistic performance estimates in stochastic environments. In contrast, reinforcement learning methods are capable of effectively adapting to these changes.

### v2

@@ 1 Introduction
2 Related Work
The Stitching Property
The concept of stitching, as discussed by (Ziebart et al., 2008), is a characteristic feature of TD-based RL algorithms, such as those described by (Mnih et al., 2013), (Lillicrap et al., 2015), (Fujimoto et al., 2018), and (Kostrikov et al., 2021), which employ Dynamic Programming techniques. This feature allows these algorithms to integrate data from different trajectories, thereby enhancing their effectiveness in managing complex tasks by leveraging historical data (Cheikhi and Russo, 2023). Recent works have shown that self-supervised learning also exhibits such property, particularly through methods like data augmentation (Shorten and Khoshgoftaar, 2019; Ghugare et al., 2024) . However, there remains a significant performance gap between these methods and algorithms that incorporate Dynamic Programming.
@@ 2 Related Work
Goal-conditioned Weighted Supervised Learning
Goal-conditioned weighted supervised learning methods (GCWSL) (Liu et al., 2021; Yang et al., 2022; Ma et al., 2022; Hejna et al., 2023) are new outstanding goal-conditioned algorithms which enable agent to reach multiple goals in GCRL. These methods repeatedly perform imitation learning on self-collected, relabeled data without any RL updates. Furthermore, these methodologies undergo theoretical validation to confirm that self-supervised learning from relabeled experiences optimizes a lower bound on the GCRL objective. This validation guarantees that the learning process is both efficient and underpinned by a rigorous theoretical foundation.
@@ 3.1 Goal-conditioned Reinforcement Learning
Q-learning
methods optimizes $\mathcal{J}(\pi)$ via the gradient computation:
@@ 3.2 Goal-conditioned Weighted Supervised Learning
where $\mathcal{B}_{r}$ denotes relabeled replay buffer, $g=\phi(s_{i})$ denotes the relabeled goals for $i\geq t$
@@
and $f(A)$ is a function about advantage $A:=A(s_{t},a_{t},g)$. The weighted function $f(A)$ exists various forms in GCWSL methods (Ghosh et al., 2021; Yang et al., 2022; Ma et al., 2022; Hejna et al., 2023). (See Section A.3 for the specific weight function meaning). Therefore GCWSL includes typical two process, acquiring sub-trajectories corresponding to state-goal (i.e, $(s,g)$) pairs and imitating them. In the process of imitation, GCWSL first train the related specific weighted function, and then extract the policy with the Equation 4.
Theorem 1.
@@
$\pi_{k+1}=\underset{\pi\in\Pi}{\operatorname*{\arg\max}}\mathbb{E}_{(s,g,s^{\prime})\sim\mathcal{B}_{r},a_{t}\sim\pi(s,g)}\left[A^{\pi_{k}}(s,a,g)\right],\text{s.t.}~{}\mathcal{D}_{\mathrm{KL}}\left(\pi\|\pi_{relabel}\right)<\epsilon,$ (5)
where $A^{\pi_{k}}=Q^{\pi_{k}}(s,a,g)-V^{\pi_{k}}(s,g)$, $\mathcal{D}_{\mathrm{KL}}$ is the KL-divergence and $g$ is the goals from relabeled data. See Section A.1 for detail proof. This theorem indicate GCWSL guarantees policy improvement when learning from the relabelled data.
@@ 4.1 Trajectory Stitching Example
Consider an example illustrated in Figure 1, where the goals $g\in\left[g^{a},g^{b}\right]$ are reachable from the states $s\in\left[s_{0}^{a},s_{0}^{b},s_{t}\right]$ although not explicitly present in the off-policy replay buffer. During training, GCWSL is updated with (state, goal) pairs sampled from the same trajectory, such as $(s_{0}^{a},g^{a})$ and $(s_{0}^{b},g^{b})$. However, during testing, when unseen (state, goal) pairs such as $(s_{0}^{a},g^{b}),(s_{0}^{b},g^{a})$ originating from distinct trajectories are introduced, GCWSL cannot find the correct actions (proof can be found in Figure 1 and Lemma 4.1 of (Ghugare et al., 2024)). In contrast, TD-based RL methods such as Q-learning can output the correct actions for $(s_{0}^{a},g^{b}),(s_{0}^{b},g^{a})$ due to the Dynamic Programming component. For instance, consider a value estimation example, $V(s_{t},g)=r+V(s_{t+1},g)$. For the pair $(s_{t+1},g)$, provided that $g$ is accessible and the value $V(s_{t+1},g)$ is accurately approximated, Q-learning can back-propagate the state-goal (state-action-goal) values to preceding pairs. An example from the off-policy replay buffer illustrates the state $s_{t}$ as a "hub state" shared by two initial trajectories, from which all goals $g^{i}\sim[g^{a},g^{b}]$ are reachable. The values of these goal-conditioned states $Q(s_{t},a_{t}^{i},g^{i})$ can be computed and recursively back-propagated to $Q(s_{0},a_{0}^{i},g^{i}),i\in\left[a,b\right]$. Q-learning has the capability to propagate rewards through paths that are connected retrospectively. This allows Q-learning to determine the optimal actions for state-goal pairs that have not been encountered together previously, such as $(s_{0}^{a},g^{b})$ and $(s_{0}^{b},g^{a})$. By iteratively back-propagating value estimates through all potential trajectory combinations, Q-learning ensures that values are consistently updated, thereby facilitating the correct action selection. In summary, Q-learning can link (state, goal) pairs across different trajectories, a capability that GCWSL lacks.
@@ 4.2 Bounding KL-Constrained Values and Q-WSL
If we he replace the advantage function $A^{\pi}(s,a,g)$ in Equation 5 by Q-function $Q^{\pi}(s,a,g)$ we have the policy improvement objective:
$\pi_{k+1}=\arg\max_{\pi}\mathbb{E}\left[Q^{\pi_{k}}(s,\pi_{k}(s,g),g)\right],\\ \;\int_{a}\pi(a|s,g)\boldsymbol{d}a=1,\;\mathbf{s.t.}~{}\operatorname{KL}(\pi\|\pi_{relabel})\leq\epsilon.$
@@
$\mathcal{L}(\lambda,\pi^{H})=\mathbb{E}_{a\sim\pi(\cdot|s,g)}\left[Q^{\pi}(s,a,g)\right]+\\ \lambda\mathbb{E}_{a,s,g\sim\mathcal{B}_{r}}\log(\pi(a|s,g)).$
For a deterministic policy $\pi$, it can be regarded as a Dirac-Delta function. Consequently, the constraint $\int_{a}\pi(a|s,g)\boldsymbol{d}a\\ =1$ is always satisfied. Therefore, Section 4.2 simplifies to:
@@
Based on this discussion and the interoperation of these two objectives, we combine Section 4.2 with Equation 5 to optimize the policy concurrently. This leads to our Q-WSL policy optimization objective for policy updates:
$\mathcal{J}_{Q-WSL}=\mathbb{E}_{(s,a,g)\sim\mathcal{B}_{r}}\big{[}\underbrace{Q^{\pi}(s,a,g}_{\text{Q-learning}})+\eta\underbrace{f(A^{\pi}(s,a,g))\cdot\log\pi_{\theta}(s,a,g)}_{\text{Advantage-weighted Regression}}\big{]},$ (8)
@@ 4.3 Practice Algorithm
We adopt the DDPG+HER as our baseline framework following (Fang et al., 2019; Yang et al., 2021), therefore for a deterministic policy $\pi(s,g)$, the Equation 16 and Equation 6 are used during training procedure. Moreover, we follow WGCSL (Yang et al., 2022) to define $f(A)$ equal to $\gamma^{i-t}\exp_{clip}(A(s_{t},a_{t},g))\cdot\epsilon(A(s_{t},a_{t},g))$, where $A(s_{t},a_{t},g)=r(s_{t},a_{t},g)+\gamma V(s_{t+1},g)-V(s_{t},g)$ and the value function $V(s_{t},g)$ is estimated using Q-value function as $V(s_{t},g)=Q(s_{t},\pi(s_{t+1},g),g)$ Because in the experiment, we found that the effect of this form of $f(A)$ is better for online goal-conditioned RL, and other forms of $f(A)$ may be more friendly for offline settings. The final overview of Q-WSL is presented in Algorithm 2.
Compared with WGCSL
Motivated by (Chane-Sane et al., 2021; Yang et al., 2022), we also compared with WGCSL and have the following boundary:
Theorem 2.
Suppose function $\exp_{clip}(A(s_{t},a_{t},g)\cdot\epsilon(A(s_{t}$ $,a_{t},g))\geq 1,g=\phi(s_{i}),i\geq 1$ over the state-action-goal combination. Consider a finite-horizon discrete Markov Decision Process (MDP), a deterministic policy $\pi$ that selects actions with non-zero probability, and a sparse reward function $r(s_{t},a_{t},g)=\mathbb{1}\left[\phi(s_{i})=g\right]$, where $\mathbb{1}\left[\phi(s_{i})=g\right]$ is an indicator function. Given trajectories $\tau=(s_{1},a_{1},\cdots,s_{T},a_{T})$ and a discount factor $\gamma\in(0,1]$, the following bounds hold:
@@
We defer the proof to Section A.1. This theorem indicates that under a reward of $r(s_{t},a_{t},g)=\mathbb{1}\left[\phi(s_{i})=g\right]$, the optimization objective of Q-WSL is at least better than that of WGCSL and GCSL.
Algorithm 2 Q-WSL for Goal-conditioned RL
@@
13:         Update $\pi_{\theta}$ with maximize goal-conditioned policy optimization with Equation 16 and Equation 6
@@
4.4 Compared with Previous Goal-conditioned Methods
@@ 5 Experiments
Benchmarks
We use some standard goal-conditioned RL research benchmarks (Plappert et al., 2018) including one task on the $Shadow-hand$ and four tasks on the $Fetch$ robot (See Figure 2).
Figure 2: Examples of goal-conditioned benchmark environments: (a) FetchReach, (b) FetchPush, (c) FetchSlide, (d) FetchPickAndPlace, (e) HandReach.
Baselines
In this section, we compare our method Q-WSL with prior works on various goal-conditioned policy learning optimization algorithms as show in Table 1 . We conducted a series of comparative experiments by implementing the baseline methods within the same framework as our proposed approach. Detailed baselines algorithm implementation is shown in Section A.3. We provide a brief overview of our baseline algorithms as bellow:
@@
Architecture Setup
We utilize the DDPG (Paul et al., 2018) in conjunction with HER as the foundational algorithm for goal-conditioned RL, against which we benchmark various baselines. Specifically, the critic architecture $Q(s,a,g)$ adheres to the Universal Value Function Approximators (UVFA) model (Schaul et al., 2015). The actor network generates actions parameterized by a Gaussian distribution. The actor and critic networks are structured with three hidden layers of multi-layer perceptrons (MLPs), each comprising 256 neurons and employing ReLU activations. The architectural setup can be denoted as [linear-ReLU]×3, followed by a linear output layer. We follow the optimal parameter settings for DDPG+HER as specified in (Plappert et al., 2018). All baseline hyperparameters are maintained consistently. Additional details are provided in Section A.5.
Evaluation Setup
For each baseline and task, we conducted evaluations using random five seeds (e.g, $\{100,200,300,400,500\}$). The policy was trained for 1000 episodes per epoch. Upon completing each training epoch, the policy’s performance was measured by calculating the average success rate from 100 independent rollouts, each using randomly selected goals. These success rates were averaged across five seeds and plotted over the learning epochs, with the standard deviation illustrated as a shaded region on the graph.
5.1 Performance Evaluation on Multi-goal Benchmark Results
@@ 5.2 Sample Efficiency
To evaluate the sample efficiency of Q-WSL compared to baseline methods, we analyze the number of training samples (i.e., $\left\langle{s,a,g^{\prime},g}\right\rangle$ tuples, $g^{\prime}$ is the relabeled goal) required to attains a specific mean success rate. For this comparison, we have selected competitive baselines that employ goal relabeling strategies in the HandReach environment. The results are depicted in Figure 3. In the HandReach environment, as shown in Figure 3, Q-WSL requires only 1000 samples to attains the 0.1 mean success rate, demonstrating superior sample efficiency. In this case, Q-WSL is at least twice the sample efficiency of previous works. The reason why a state-of-the-art baseline DWSL is not listed here is that there is basically no success rate in the HandReach environment. Overall, Q-WSL enhances sample efficiency by an average factor of seven (6.8) compared to the baseline methods.
@@ 5.3 Robust to Environmental Stochasticity
As we see in Figure 4, Q-WSL is the most robust to stochasticity in the FetchPush environment, also outperforming baseline algorithms in terms of mean success rate under various noise levels. WGCSL exhibits minimal sensitivity to variations in noise levels, whereas DDPG+HER is more sensitive, and GoFar is the most affected by noise fluctuations. Despite DWSL’s insensitivity to noise, its overall performance remains suboptimal. We analyze that the assumption of deterministic dynamics in the self-supervised learning methods such as WGCSL, GoFar, and DWSL may lead to overly optimistic performance estimates in stochastic environments. In contrast, reinforcement learning methods are capable of effectively adapting to these changes.